OpenStudy (anonymous):

write the equation of the line in slope-intercept form that is parallel to the given line and passes through the given point. y=8x+5 , (6,7)

OpenStudy (yanasidlinskiy):

Is (6,7) your answer? Or what?

OpenStudy (anonymous):

thats part of the problem

OpenStudy (rea201):

Since the line in parallel to the line equation given the new line will have the same slope as the one give

OpenStudy (anonymous):

No, (6,7) is not the answer. I usually just graph this..

OpenStudy (rea201):

Now since we now thw slope is 8 we can form the equation using the point given

OpenStudy (anonymous):

@Kaylaaa Do you know what slope-intercept form is?

OpenStudy (rea201):

(y-y1)=m(x-x1)

OpenStudy (rea201):

Do you know where we go from here @Kaylaaa

OpenStudy (anonymous):

Slope-intercept form is y - mx + b.

OpenStudy (anonymous):

* y = mx + b

OpenStudy (anonymous):

this is all just so confusing to me

OpenStudy (anonymous):

If you want a PARALLEL line you MUST have the SAME slope.

OpenStudy (rea201):

But @greenlegodude57 I am showing her how to do it analytically

OpenStudy (anonymous):

The slope is the 'mx' part.

OpenStudy (rea201):

I am teaching her so let me finish first

OpenStudy (anonymous):

The slope in the first equation is 8, so you know that your second equation must have 8 for its slope also(So it can be parallel with the first equation).

OpenStudy (anonymous):

And your second equation must cross through the point (6, 7)

OpenStudy (rea201):

Now @Kaylaaa we are going to be using the equation i gave. m=8 y1=7 and x1=6

OpenStudy (anonymous):

The only thing that's going to change is the number at the end of the first equation to create your second equation.

OpenStudy (rea201):

so we get y-7=8(x-6)

OpenStudy (rea201):

now we can change it to slope intercept form

OpenStudy (anonymous):

y = 8x - 33 will be your answer.

OpenStudy (anonymous):

ohkay ..

OpenStudy (anonymous):

wait which one is it ?

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

Oops my bad. It's y = 8x - 41

OpenStudy (rea201):

ignore what i just said what i origanally put is correct sorry for the confusion. 6times 8 is 48

OpenStudy (anonymous):

y = 8x - 33 crosses through (5, 7) not (6, 7). So your answer is y = 8x - 41

OpenStudy (rea201):

the answer is y=8x-41

OpenStudy (rea201):

i graphed it to check my answer and it is absolutely correct

OpenStudy (anonymous):

well thank you .! (: